# Enhancing Redundancy-based Automated Program Repair by Fine-grained Pattern Mining

### v1

@@ 4.2. Configuration and Metrics
For each bug, Repatt first constructs the token-level code patterns based on the complete faulty program under repair before the online repair (see Section 3.1). The construction is actually efficient and on average took about three minutes in our experiment. Then, given the faulty line of code, Repatt first generates at most 200 patches based on the constructed token-level code patterns and then generates at most 1000 patches based on the expression-level code patterns due to its large search space. Finally, all the patches will be ranked based on the ranking strategy introduced in Section 3.4. Additionally, we assign each bug 90 minutes for repair (Le et al., 2016; Wen et al., 2018; Saha et al., 2017). We empirically set the default value of MIN_SUPPORT in token-level pattern mining as 3, and the MAX_SKIP and MAX_LEN as 8 and 5, respectively. For baselines, we adopt their published experimental results directly in their open-source repositories. In particular, for the three large-scale pre-trained deep learning-based methods, we adopt the fine-tuned models CodeT5-large, CodeGen-6B, and PLBART-large, which achieved the best repair results (Jiang et al., 2023a).
@@
In our experiment, Repatt generates at most 3 candidate patches that can pass all the test cases (i.e., plausible patches) for each bug since some baseline approaches did not report the ranking of patches, e.g., SelfRepair. Following previous study (Xin and Reiss, 2017b; Jiang et al., 2018; Liu et al., 2019; Zhu et al., 2021), a patch is deemed to be correct iff it is semantically-equivalent to the developer patch by manual check. Finally, we report the number of correctly repaired bugs (equivalent to the well-known recall) and the precision of patches (the ratio of correctly repaired bugs to all the bugs that have plausible patches). Please note the recall and precision here are different from those in the deep learning field since all the bugs are expected to be repaired, i.e., the true negative in deep learning field does not make sense here. All experiments were conducted on a server with Ubuntu 18.04, equipped with 128GB RAM and a processor of Intel(R) Xeon(R) E5-2640.
5. Result Analysis
@@
Table 1. Number of bugs repaired by different methods with perfect fault localization.
@@
Project Recoder TBar SimFix TransplantFix SelfRepair CodeT5 CodeGen PLBART Repatt Combine
Defects4J v1.2 (395 bugs) Mockito 2/- 3/3 0/0 3/3 3/3 4/5 3/4 2/3 1/2 2/5
Closure 23/- 16/24 5/6 10/19 19/23 6/9 9/12 4/10 10/12 20/31
Chart 10/- 11/13 4/8 6/10 7/10 4/4 7/7 4/4 6/7 11/21
Lang 10/- 13/18 8/13 4/10 10/13 6/8 5/5 4/4 6/11 16/27
Math 18/- 22/35 14/26 12/25 21/25 12/14 12/14 13/16 15/25 22/51
Time 3/- 3/6 1/1 1/2 3/3 1/1 2/2 2/2 2/3 4/7
Defects4J v2.0 (440 bugs) Closure 0/- 0/0 1/1 0/3 1/1 1/1 1/1 1/1 1/1 2/5
Cli 3/- 1/7 0/1 4/8 8/9 2/3 4/5 2/3 4/9 4/15
Codec 2/- 3/6 1/1 2/4 8/9 1/1 2/3 1/1 4/6 3/8
Collections 0/- 0/0 0/0 0/0 1/1 0/0 1/1 0/0 0/0 0/0
Compress 3/- 2/12 0/4 4/10 6/8 3/5 3/6 3/5 4/13 7/18
Csv 3/- 2/6 0/2 1/2 1/1 1/2 2/3 2/4 1/3 2/7
Gson 0/- 1/4 2/2 1/2 1/1 0/0 0/0 0/0 2/3 2/5
JacksonCore 0/- 0/5 0/2 1/6 3/3 3/3 2/3 2/2 3/7 3/9
JacksonDatabind 0/- 2/17 1/10 7/20 8/10 4/7 3/6 2/4 10/18 11/33
JacksonXml - 0/1 0/0 0/0 1/1 0/0 0/0 0/0 0/0 1/1
Jsoup 7/- 6/17 1/2 3/8 6/8 4/6 5/7 4/5 5/13 10/25
JxPath 0/- 1/6 0/0 2/7 1/1 0/0 0/1 0/1 1/5 4/12
Total 84/- 85/180 38/79 61/139 108/130 52/ 69 61/80 46/ 65 75/138 124/280
Precision(%) - 47.2 48.1 43.9 83.1 75.4 76.3 70.8 54.3 44.3
5.1. Overall Effectiveness of Repatt (RQ1)
5.1.1. Perfect fault localization
As explained in Section 4.1, we evaluated the effectiveness of our approach by comparing it with eight state-of-the-art APR approaches. The repair results are presented in Table 1 when given the actual faulty locations by following existing studies (Jiang et al., 2023a; Yang et al., 2023; Zhu et al., 2021; Liu et al., 2019; Ye et al., 2022b). In each cell, X/Y denotes the repair method can generate at least one plausible patch (i.e., can pass all the test cases) for Y bugs on different projects, where X bugs are indeed repaired correctly. From the table we can see that our approach Repatt successfully repaired 75 bugs but also generated incorrect patches for other 63 bugs, leaving the patch precision 54.3%. In particular, 66 bugs were correctly repaired by the first plausible patch. As a result, developers may only consider the first patch generated by Repatt for manual validation to save human effort in practice. The results also show that our approach outperforms the state-of-the-art redundancy-based methods SimFix and TransplantFix by repairing 97.4% and 23.0% more bugs, indicating the effectiveness of our approach in reusing code at different granularities for patch generation. Please note that SelfRepair repaired 108 bugs when considering Top-50 patches. When considering the first (i.e., Top-1) plausible patches, it can only repair 34 bugs as reported in the original paper (Ye et al., 2022a). Please note that the latest advanced learning-based repair tool KNOD (Jiang et al., 2023b) did not report the detailed experimental results, i.e., the number of bugs that were correctly repaired by Top-1 patches. Nevertheless, even considering Top-10 patches, KNOD could only repair 20 and 13 bugs from Defects4J v1.2 and v2.0, respectively, still underperfoming our approach. Compared with the others, our approach can also achieve similar effectiveness. For example, Repatt achieved higher patch precision than TBar although repaired fewer bugs. In contrast, Repatt repaired more bugs than the pre-trained models (CodeT5, CodeGen, and PLBART) but achieved a lower patch precision. Moreover, as it will be presented in Section 5.1.2, our approach complements existing APRs by repairing many new bugs. Nevertheless, although our approach can effectively repair many bugs, there are still many bugs that cannot be repaired by Repatt because the patches cannot be assembled from individual reference code. To further improve the correct repair, enlarging the patch space may help, such as combining multiple reference code or searching a code base beyond the faulty project for patch generation. We plan to further investigate its performance in the near future.
5.1.2. Degree of complementary
According to the results shown in Table 1, although Repatt can successfully repair many bugs, it is unknown whether Repatt is still valuable in the face of many state-of-the-art APR approaches. Therefore, we further investigated the extent to which our approach complements the state-of-the-art baselines. We analyzed the overlaps of bugs repaired by both Repatt and the baselines. Specifically, we classified the baselines into traditional and deep-learning methods according to the patch generation techniques, and then compared Repatt with these two methods separately. Figure 5 presents the results. In summary, Repatt successfully repaired 19 unique bugs that the traditional APRs cannot fix, and 15 unique bugs that the deep learning methods cannot fix. In particular, when compared with all the eight baselines, our approach still can repair 9 unique bugs. For example, to repair the bug shown in Listing 1, the constant value “4” should be replaced by “3”. Without the guidance of the fine-grained reference code, it is nearly impossible for existing APRs to get the correct patch due to the large search space. In general, compared with traditional APRs, Repatt can more effectively utilize fine-grained reusable code elements for patch generation, while compared with deep learning methods, Repatt only depends on a small number of reusable samples rather than the large-scale training data, which enables Repatt to repair infrequent bugs. In conclusion, the results reflect that our approach indeed complements existing approaches.
@@ 5.1.2. Degree of complementary
(b) DL methods
Figure 5. Overlaps of bugs repaired by different approaches.
5.1.3. Automated fault localization
@@
In addition, some of the baseline approaches (i.e., SimFix, TBar, TransplantFix, and Recoder.) were also evaluated in the scenario where the perfect fault localization was unknown in the original paper. Therefore, we also compared our approach with them when using the automated fault localization results by following existing research (Liu et al., 2019; Jiang et al., 2018; Zhu et al., 2021). Specifically, we utilized the commonly-used Ochiai(Abreu et al., 2007) algorithm that is implemented by the GZoltar toolset (Campos et al., 2012) to obtain a list of candidate faulty locations like existing APR tools (Liu et al., 2019; Jiang et al., 2018; Zhu et al., 2021). In our experiment, we use the default configuration of GZoltar. Table 2 summarizes the experimental results. From the table, we can see that the number of correctly repaired bugs drops sharply for all the repair approaches. For example, SimFix, TBar, TransplantFix, Recoder, and Repatt repaired 5.2%, 41.2%, 27.9%, 20% and 60% fewer bugs, respectively, compared with providing the perfect fault localization. The large decline in our approach is mainly due to the inaccurate fault localization results, which will cause a large search space since the finer-grained code reuse tends to generate candidate patches, making the correct patches unfound within the given time budget. Moreover, the patch precision of all the baseline approaches also incurred a large decline. For instance, the precision of SimFix, TBar, and TransplantFix drops 7.8%, 19.1%, and 26.9%, respectively. On the contrary, the precision of our approach was improved by 54.3%. The reason is that fine-grained code changes at the faulty location directly affect the results of failing tests and thus have a higher possibility to make them pass without affecting the results of passing tests, while fine-grained code changes to the correct code will have a larger impact on the passing tests. In contrast, the coarse-grained code changes by existing APRs tend to affect both failed and passed tests, and thus they will generate incorrect patches more easily under inaccurate fault localization due to the problem of weak tests (Qi et al., 2015; Xiong et al., 2018; Xin and Reiss, 2017a). Notice that the high precision of patches is very important since it affects the usability of APR techniques (Qi et al., 2015; Xiong et al., 2018), especially in the real-world repair scenarios where the prefect fault localization is unknown – high precision denotes less wasted human effort for manual validation.
@@
Table 2. Repair result comparison when using SBFL
@@
SimFix TBar TransplantFix Recoder Repatt Combine
Total 36/81 50/131 44/137 70/140 31/37 90/236
Precision (%) 44.4 38.2 32.1 50.0 83.8 38.1
5.2. Improvement over SOTA APRs (RQ2)
@@
As presented in Section 5.1.2, Repatt complements existing APR approaches. In this research question, we aim to explore the possibility of our approach to improving existing APRs. Specifically, we combine Repatt with existing APRs by proposing a post patch ranking strategy and see whether it can further improve the best-performing APR. Given the patches generated by different APRs for a bug, we will rank the ones with smaller code changes higher by following the insights from both our study and the previous research (Mechtaev et al., 2015). In particular, we use GumTree (Falleri et al., 2014), a fine-grained and mature code differencing tool, to measure the change sizes of patches. If two patches have the same number of code changes, we use the patch precision to break the tie. In this study, we combined our approach with the three traditional APR approaches, i.e., Simfix, TBar, and TransplantFix. In this way, the combined method shares a similar repair pipeline to the individuals, i.e., not demanding large-scale training data. Therefore, the patch ranking will be Repatt$>$SimFix$>$TBar$>$TransplantFix based on their patch precision in Tables 1 and 2 for patches having the same change size.
@@
Figure 6. Source of correct patches in Combine
@@
The experimental results are also presented in Tables 1 and 2 (i.e., Combine) when using the perfect fault localization and the SBFL, respectively. The results show that by combining the strength of different APR tools, Combine correctly repaired 124 bugs by the first patches, 39 more bugs compared with the best-performing APR (TBar for Top-1 patches), leaving the improvement as 45.9%. Similarly, when using the SBFL results, Combine also outperforms the best-performing Recoder by repairing 20 more bugs, and improves the individuals by at least 80% (vs Tbar). Our results show the promise to improve the repair capability of APRs by combining the strength of individuals. In particular, to analyze the contribution of our approach in the combination, we present the percentages of correct patches generated by each APR in Figure 6. About 41.1% and 25.6% correct patches were contributed by Repatt under different repair settings, indicating its large contribution to the overall effectiveness of the combined method.
However, though the combined method is effective in repairing much more bugs, the patch precision is still low, i.e., less than 45%. To improve the patch precision, existing patch filtering approaches (Tan et al., 2016; Xiong et al., 2018; Xin and Reiss, 2017a) can be further incorporated. As far as we are aware, this is the first attempt to combine the strength of different APRs. Although simple, the combined method already shows its promise. In the future, we plan to explore more effective combining methods and also incorporate the patch filtering process for improving the patch precision. Nevertheless, the results obtained from this study encourage further research in this direction.
5.3. Contribution of Each Component (RQ3)
@@
Repatt incorporates two major components for patch generation, i.e., offline token-level pattern mining and expression-level code search. To evaluate the effectiveness of each component, we have repeated our experiment by removing each component, respectively. Table 3 presents the experimental results. The offline token-level pattern mining contributed 42 correct fixes and expression-level code search contributed 33 correct fixes, demonstrating that both components largely contributed to the overall effectiveness of Repatt. However, from the table, we can also observe that the expression-level code search tends to generate more plausible but incorrect patches compared with the token-level pattern mining, which is consistent with the conclusions in prior studies (Mechtaev et al., 2015). Our further analysis of the results shows that most of incorrect patches are generated due to referencing infrequent code. In contrast, token-level pattern mining takes the frequency into consideration, which potentially can effectively filter out many incorrect patches. In summary, the two components in Repatt complement each other.
@@
In addition, as introduced in Section 3, our token-level pattern mining process incorporates a skip-fashion pattern mining process, which can find more potential reference patterns for patch generation. Therefore, we further conducted an experiment, where we only use the token-level repair but disable the skip-fashion in pattern mining (i.e., MAX_SKIP=0 in Algorithm 1). The column of “No Skip” in Table 3 presents the results. It shows that this skip-fashion mining process is effective since its removal caused 42-24=18 fewer bugs to be repaired. Moreover, the patch precision is also decreased from 65.6% to 44.4%, demonstrating the value of this process in Repatt. Additionally, we further evaluated the performance of the S-TAC form in our expression-level repair. Specifically, we did not transform the reference code into S-TAC. Instead, they will be decomposed into normal expressions without removing the code structures based on the abstract syntax tree. The experimental results are also presented in Table 3 (i.e., “No S-TAC”). After removing the S-TAC form, Repatt failed to repair any bugs by searching reference code online. We further analyzed the results and found that the contexts of most reference code from online search are different from the faulty code, making the code adaptation fail to generate correct patches. On the contrary, 22 incorrect patches were generated. In summary, the experimental results demonstrate that both the skip-fashion online pattern mining and the S-TAC form in the online search process are effective.
Table 3. Number of bugs fixed by each component.
@@
5.4. The Impact of Pattern Frequency (RQ4)
As introduced in Section 4.2, the threshold of pattern frequency is 3 by default. To investigate its impact on the effectiveness of our approach, we have conducted an additional experiment when using different frequency settings for MIN_SUPPORT, i.e., 2, 3, and 5. As a consequence, when adopting different settings, the performance of Repatt was not largely affected. Specifically, Repatt correctly repaired 43/91, 42/64, and 32/68 bugs by the token-level patterns when using the threshold as 2, 3, and 5, respectively. In other words, our approach is not very sensitive regarding the number of correct patches when the threshold is small (i.e., 2 or 3). However, when the threshold is too large (such as 5), Repatt tends to repair much fewer bugs. The reason is also clear as the large threshold will cause fewer patterns that can be used for patch generation. On the contrary, a smaller threshold (i.e., 2) tends to produce much more incorrect patches (i.e., 91-43=48) since the referenced patterns may be not general enough for reuse, which will affect the usability of Repatt. In summary, the recall and precision of program repair indeed contradict each other. Effective patch generation techniques are always in an urgent need. In this paper, we have moved forward towards this direction. In general, setting the threshold as 3 can produce relatively better results.
6. Discussion

### v3

@@ IV-B Configuration and Metrics
For each bug, Repatt first constructs the token-level code patterns based on the complete faulty program under repair before the online repair (see Section III-A). The construction is actually efficient and on average took about three minutes in our experiment. Then, given the faulty line of code, Repatt first generates at most 200 patches based on the constructed token-level code patterns and then generates at most 1000 patches based on the expression-level code patterns due to its large search space. Finally, all the patches will be ranked based on the ranking strategy introduced in Section III-D. For baselines, we adopt their published experimental results directly in their open-source repositories. Specifically, as the original CapGen does not provide experimental results on the Closure and Mockito bugs from Defects4J [41], we supplement these results using those results reported in a previous study [54]
@@
In our experiment, Repatt generates at most 3 candidate patches that can pass all the test cases (i.e., plausible patches) for each bug since some baseline approaches did not report the ranking of patches, e.g., SelfAPR and LLM-based methods. Following previous study [9, 8, 4, 18], a patch is deemed to be correct iff it is semantically equivalent to the developer patch by manual check. In this process, the first three authors independently conducted the annotation and reached a consensus through discussion. We also published all our results for further inspection and verification. Finally, we report the number of correctly repaired bugs (equivalent to the well-known recall) and the precision of patches (the ratio of correctly repaired bugs to all the bugs that have plausible patches).
All experiments were conducted on a server with Ubuntu 18.04, equipped with 128GB RAM and a processor of Intel(R) Xeon(R) E5-2640.
V Result Analysis
@@
TABLE I: Number of bugs repaired by different methods with perfect fault localization. In the table, x/y denotes the corresponding approach generates correct patches for x bugs and generates plausible patches for y bugs.
@@
Project TBar SimFix TransplantFix SelfAPR Recoder AlphaRepair Repilot GAMMA Repatt Combine
Defects4J v1.2
(395 bugs)
Mockito 3/3 0/0 3/3 3/3 2/- 4/- 0/- 2/- 1/2 2/5
Closure 16/24 5/6 10/19 19/23 23/- 22/- 21/- 23/- 10/12 20/31
Chart 11/13 4/8 6/10 7/10 10/- 8/- 5/- 10/- 6/7 11/21
Lang 13/18 8/13 4/10 10/13 10/- 12/- 14/- 15/- 6/11 16/27
Math 22/35 14/26 12/25 21/25 18/- 20/- 20/- 24/- 15/25 22/51
Time 3/6 1/1 1/2 3/3 3/- 2/- 1/- 2/- 2/3 4/7
Defects4J v2.0
(440 bugs)
Closure 0/0 1/1 0/3 1/1 0/- 0/- 0/- 0/- 1/1 2/5
Cli 1/7 0/1 4/8 8/9 3/- 5/- 6/- 9/- 4/9 4/15
Codec 3/6 1/1 2/4 8/9 2/- 6/- 6/- 3/- 4/6 3/8
Collections 0/0 0/0 0/0 1/1 0/- 0/- 1/- 0/- 0/0 0/0
Compress 2/12 0/4 4/10 6/8 3/- 1/- 3/- 4/- 4/13 7/18
Csv 2/6 0/2 1/2 1/1 3/- 1/- 3/- 0/- 1/3 2/7
Gson 1/4 2/2 1/2 1/1 0/- 2/- 1/- 3/- 2/3 2/5
JacksonCore 0/5 0/2 1/6 3/3 0/- 3/- 3/- 3/- 3/7 3/9
JacksonDatabind 2/17 1/10 7/20 8/10 0/- 8/- 8/- 10/- 10/18 11/33
JacksonXml 0/1 0/0 0/0 1/1 0/- 0/- 0/- 0/- 0/0 1/1
Jsoup 6/17 1/2 3/8 6/8 7/- 9/- 18/- 11/- 5/13 10/25
JxPath 1/6 0/0 2/7 1/1 0/- 1/- 1/- 2/- 1/5 4/12
Total 85/180 38/79 61/139 108/130 84/- 104/- 111/- 121/- 75/138 124/280
Precision(%) 47.2 48.1 43.9 83.1 - - - - 54.3 44.3
V-A Overall Effectiveness of Repatt (RQ1)
V-A1 Perfect fault localization
As explained in Section IV-A, we evaluated the effectiveness of our approach by comparing it with eight state-of-the-art APR approaches since ITER was only evaluated under the automated fault localization setting (ref. Section V-A3). The repair results are presented in Table I when given the actual faulty locations by following existing studies [20, 33, 18, 4, 19]. From the table we can see that Repatt successfully repaired 75 bugs while generating incorrect patches for other 63 bugs, yielding a patch precision of 54.3%. Notably, 66 bugs were correctly repaired by the first plausible patch, suggesting that developers could focus primarily on the first patch generated by Repatt to reduce manual validation efforts. The results also show that our approach outperforms the state-of-the-art redundancy-based methods SimFix and TransplantFix by repairing 97.4% and 23.0% more bugs, indicating the effectiveness of our approach in reusing code at different granularities for patch generation. Compared to SelfAPR, which repaired 108 bugs when considering Top-50 generated patches but only 34 when restricted to Top-1, Repatt shows a more favorable trade-off between precision and recall. When analyzing the incorrect patches generated by Repatt, we found that most of them resulted from the identification of non-representative patterns, which led to incorrect modification locations or inappropriate fixes. However, due to the weakness of the test suites [48, 55, 46], these patches were able to pass all test cases, highlighting a common challenge in test-based APR evaluation.
Similarly, LLM-based repair methods also did not report the ranking of their correct patches. Specifically, AlphaRepair [22] and Repilot [23] generated up to 5,000 patches per bug, while GAMMA [24] even did not limit the number of candidate patches. Although these methods may achieve a higher number of correct fixes, they require developers to spend significant time manually reviewing patches, reducing their practicality [46, 48]. In contrast, Repatt enforces strict limits on the number of generated patches, effectively mitigating this issue. Furthermore, we analyzed Repatt’s performance on repairing multi-location bugs. Among 75 correctly repaired bugs, 30 involved multiple locations, highlighting the effectiveness of Repatt’s fine-grained repair strategy.
@@ V-A1 Perfect fault localization
(b) LLM and DL based methods
(c) All methods
Figure 5: Overlaps of bugs repaired by different approaches.
V-A2 Degree of complementary
We also analyzed the overlaps of bugs repaired by both Repatt and the baselines. Specifically, we classified the baselines into traditional, LLM-based and deep-learning-based methods according to the patch generation techniques, and then compared Repatt with these two methods separately. Figure 5 presents the results. In summary, Repatt successfully repaired 19 unique bugs that the traditional APRs cannot fix, and 9 unique bugs that the LLM-basd and deep learning-based methods cannot fix. In particular, when compared with all the eight baselines, our approach still can repair 5 unique bugs. For example, to repair the bug shown in Listing 1, the constant value “4” should be replaced by “3”. Without the guidance of the fine-grained reference code, it is nearly impossible for existing APRs to get the correct patch due to the large search space. In general, compared with traditional APRs, Repatt can more effectively utilize fine-grained reusable code elements for patch generation. To better understand its repair capabilities, we manually analyzed the 19 bugs that were uniquely fixed by Repatt. We found that 7 of them (36.8%) involved fine-grained code reuse and edits, such as changing a single character, a variable name or a method name. This demonstrates the capability of Repatt to precisely identify and apply fine-grained code changes effectively, which traditional approaches may overlook due to their coarse-grained search spaces. Moreover, while compared with LLM-based and deep learning-based methods, Repatt only depends on a small number of reusable samples rather than the large-scale training data, which enables Repatt to repair infrequent bugs, especially those requiring domain-specific knowledge. In conclusion, the results reflect that our approach indeed complements existing approaches.
V-A3 Automated fault localization
@@
In addition, some of the baseline approaches [8, 4, 18, 33, 53] were also evaluated in a more realistic scenario where the perfect fault localization was unknown. To ensure a fair comparison, we evaluated our approach under the same conditions. Specifically, we utilized the commonly-used Ochiai [56] algorithm, implemented by the GZoltar toolset [57], to obtain a list of candidate faulty locations like existing APR tools [4, 8, 18]. Table II summarizes the experimental results. From the table, we can see that the number of correctly repaired bugs drops sharply for all the repair approaches due to the inaccurate fault localization results. Moreover, almost all baseline approaches experienced a significant decline in patch precision. For instance, the precision of SimFix, TBar and TransplantFix drops 7.8%, 19.1%, and 26.9%, respectively. In contrast, the precision of our approach was improved by 54.3%. One of the reasons is that fine-grained code changes focus on fixing local code without making extensive changes to the original content, thereby preserving program logic and maintaining the effectiveness of test cases, leading to more correct patches. In contrast, coarse-grained modifications often involve larger code changes, which may unintentionally alter program functionality, resulting in many patches that pass the tests but are semantically incorrect due to the problem of weak tests [46, 48, 47]. In summary, our approach significantly outperforms the baselines by achieving 15.6%-51.7% higher patch precision. Notice that the high precision of patches is very important since it affects the usability of APR techniques [46, 48], especially in the real-world repair scenarios where perfect fault localization is unknown – high precision denotes less wasted human effort for manual validation.
@@
TABLE II: Repair result when using SBFL (#correct/#plausible).
@@
CapGen SimFix TBar TransplantFix Recoder ITER Repatt Combine
Total 22/49 36/81 50/131 44/137 70/140 45/66 31/37 90/236
Precision (%) 44.90 44.4 38.2 32.1 50.0 68.2 83.8 38.1
V-B Improvement over SOTA APRs (RQ2)
@@
As presented in Section V-A2, Repatt complements existing APR approaches. In this research question, we aim to explore the possibility of our approach to improving existing APRs. Specifically, we combine Repatt with existing APRs by proposing a post patch ranking strategy and see whether it can further improve the best-performing APR. Given the patches generated by different APRs for a bug, we will rank the ones with smaller code changes higher by following the insights from both our study and the previous research [51]. In particular, we use GumTree [58], a fine-grained and mature code differencing tool, to measure the change sizes of patches. If two patches have the same number of code changes, we use the patch precision to break the tie. In this study, we combined our approach with the three traditional APR approaches, i.e., Simfix, TBar, and TransplantFix. In this way, the combined method shares a similar repair pipeline to the individuals, i.e., not demanding large-scale training data. Therefore, the patch ranking will be Repatt>>SimFix>>TBar>>TransplantFix based on their patch precision in Tables I and II for patches having the same change size.
@@
Figure 6: Source of correct patches in Combine
@@
The experimental results are also presented in Tables I and II (i.e., Combine) when using the perfect fault localization and the SBFL, respectively. The results show that by combining the strength of different APR tools, Combine correctly repaired 124 bugs by the first patches, 39 more bugs compared with the best-performing APR (TBar for Top-1 patches), leaving the improvement as 45.9%. Moreover, even compared to LLM-based methods, which do not report detailed patch rankings and impose loose constraints on the number of generated patches (e.g., producing 5,000 or more patches per bug), Combine still outperforms these 3 LLM-based approaches, repairing up to 20 more bugs. Similarly, when using the SBFL results, Combine also outperforms the best-performing Recoder by repairing 20 more bugs, and improves the individuals by at least 80% (vs Tbar). Our results show the promise to improve the repair capability of APRs by combining the strength of individuals. Although this approach may increase the time required to repair one defect, the additional overhead can be mitigated through parallel execution, similar to some LLM-based APR methods [59]. In particular, to analyze the contribution of our approach in the combination, we present the percentages of correct patches generated by each APR in Figure 6. About 41.1% and 25.6% correct patches were contributed by Repatt under different repair settings, indicating its large contribution to the overall effectiveness of the combined method. However, though the combined method is effective in repairing much more bugs, the patch precision is still low, i.e., less than 45%. To improve the patch precision, existing patch filtering approaches [5, 48, 47] can be further incorporated.
V-C Contribution of Each Component (RQ3)
@@
Repatt incorporates two major components for patch generation, i.e., offline token-level pattern mining and online expression-level code search. To evaluate the effectiveness of each component, we have repeated our experiment by removing each component, respectively. Table III presents the experimental results. The offline token-level pattern mining contributed 42 correct fixes and expression-level code search contributed 33 correct fixes, demonstrating that both components largely contributed to the overall effectiveness of Repatt. However, from the table, we can also observe that the expression-level code search tends to generate more plausible but incorrect patches compared with the token-level pattern mining, which is consistent with the conclusions in prior studies [51]. Our further analysis of the results shows that most of incorrect patches are generated due to referencing infrequent code. In contrast, token-level pattern mining takes the frequency into consideration, which potentially can effectively filter out many incorrect patches. In summary, the two components in Repatt complement each other.
@@
In addition, as introduced in Section III, our token-level pattern mining process incorporates a skip-fashion pattern mining process, which can find more potential reference patterns for patch generation. Therefore, we further conducted an experiment, where we only use the token-level repair but disable the skip-fashion in pattern mining (i.e., MAX_SKIP=0 in Algorithm 1). The column of “No Skip” in Table III presents the results. It shows that this skip-fashion mining process is effective since its removal caused 42-24=18 fewer bugs to be repaired. Moreover, the patch precision is also decreased from 65.6% to 44.4%, demonstrating the value of this process in Repatt. Additionally, we further evaluated the performance of the S-TAC form in our expression-level repair. Specifically, we did not transform the reference code into S-TAC. Instead, they will be decomposed into the traditional TAC format without removing the code structures based on the abstract syntax tree. The experimental results are also presented in Table III (i.e., “No S-TAC”). After removing the S-TAC form, Repatt failed to repair any bugs by searching reference code online. We further analyzed the results and found that the contexts of most reference code from online search are different from the faulty code, making the code adaptation fail to generate correct patches. On the contrary, 22 incorrect patches were generated. In summary, the experimental results demonstrate that both the skip-fashion online pattern mining and the S-TAC form in the online search process are effective.
TABLE III: Bugs fixed by each component (#correct/#plausible).
@@
V-D The Impact of Pattern Frequency (RQ4)
As introduced in Section IV-B, the threshold of pattern frequency is 3 by default. To investigate its impact on the effectiveness of our approach, we have conducted an additional experiment when using different frequency settings for MIN_SUPPORT, i.e., 2, 3, and 5. As a consequence, when adopting different settings, the performance of Repatt was not largely affected. Specifically, Repatt correctly repaired 43/91, 42/64, and 32/68 bugs by the token-level patterns when using the threshold as 2, 3, and 5, respectively. In other words, our approach is not very sensitive regarding the number of correct patches when the threshold is small (i.e., 2 or 3). However, when the threshold is too large (such as 5), Repatt tends to repair much fewer bugs. The reason is evident as the large threshold will cause fewer patterns that can be used for patch generation. On the contrary, a smaller threshold (i.e., 2) tends to produce much more incorrect patches (i.e., 91-43=48) since the referenced patterns may be not general enough for reuse, which will affect the usability of Repatt. In summary, the recall and precision of program repair indeed contradict each other. Effective patch generation techniques are always in an urgent need. In this paper, we have moved forward towards this direction. In general, setting the threshold as 3 can produce relatively better results.
VI Discussion